# Uncertainty Quantification Metrics for Deep Regression

### v2

@@ Abstract
When deploying deep neural networks on robots or other physical systems, the learned model should reliably quantify predictive uncertainty. A reliable uncertainty allows downstream modules to reason about the safety of its actions. In this work, we address metrics for evaluating such an uncertainty. Specifically, we focus on regression tasks, and investigate Area Under Sparsification Error (AUSE), Calibration Error, Spearman’s Rank Correlation, and Negative Log-Likelihood (NLL). Using synthetic regression datasets, we look into how those metrics behave under four typical types of uncertainty, their stability regarding the size of the test set, and reveal their strengths and weaknesses. Our results indicate that Calibration Error is the most stable and interpretable metric, but AUSE and NLL also have their respective use cases. We discourage the usage of Spearman’s Rank Correlation for evaluating uncertainties and recommend replacing it with AUSE.
I Introduction
@@
In recent years, there has been a rapid advance in the adoption of neural network-based methods for many tasks in robotics. Following this adoption, increasing scrutiny has been directed towards neural network-based methods for their lack of reliability and interpretability. While neural networks have achieved impressive performance for many different tasks, the fact remains that they can be unreliable in real-world deployment [1]. Additionally, their lack of interpretability makes it difficult to know how and when they may perform unreliably. For these reasons, increasing attention has been directed at the uncertainty output from neural networks, and the importance of introspective qualities [1]. Arguably the most important introspective quality is a reliable uncertainty estimate.
Despite increasing attention being directed toward uncertainty quantification (UQ), most of the work is directed toward uncertainty in classification tasks. In robotics, regression problems are common, and there is a lack of common understanding surrounding the available metrics. In this work, we have identified four metrics that are commonly used to measure various qualities in a predicted regression uncertainty. Specifically, we investigate the Area Under Sparsification Error (AUSE) [2], Spearman’s Rank Correlation [3], Negative Log Likelihood (NLL)[4], and Calibration Error [5]. These UQ metrics measure different aspects of uncertainty that are all orthogonal to the regression task performance (typically mean squared error (MSE)) as indicated in Fig. 1.
@@
Figure 1: An illustration for UQ metrics and regression metrics. Note: the axes of calibration error and AUSE are distinct, but not orthogonal.
@@
With the help of synthetic datasets, we explore the usage of these metrics with the aim of understanding exactly what they measure, their limitations and whether or not their measurements are useful for practical applications.
@@
  • We create simple synthetic datasets that highlight different types of uncertainty: 1) Homoscedastic aleatoric uncertainty; 2) Heteroscedastic aleatoric uncertainty; 3) Multimodal aleatoric uncertainty; 4) epistemic uncertainty.
@@
  • Using our datasets, we explore how different metrics behave under different types of uncertainty.
  • We investigate how each metric behaves given the ground truth data distribution.
  • Finally, we reason about the strengths and limitations of each metric.
@@
As an additional contribution, we express a mathematical formulation for the AUSE metric. To the best of our knowledge, AUSE has previously only been described informally, using natural language.
II Related Work
@@
Historically, methods that deal with any type of regression uncertainty have been formulated using parameterized distributions. Classic examples include Kalman filters [6], hidden Markov models [7], and Gaussian processes [8]. All of these employ parameterized distributions with well-defined measures of uncertainty. Modern machine learning methods are built upon a solid statistical foundation and most methods output some form of uncertainty [9]. However, it has been observed that there is a discrepancy between the uncertainty output from modern neural networks, and the observed empirical accuracy [10]. Additionally, lacking interpretability in these uncertainties have led to the development of methods to formulate more statistically grounded uncertainties [11].
@@
Along with the research on uncertainty estimation, a number of methods have been developed to assess how trustworthy a predicted uncertainty is. Perhaps the most commonly used (for example in [10, 12, 13, 14]) is the expected calibration error (ECE) [5]. While not as common as ECE, NLL (negative log likelihood) is also used, for example in [15, 16, 17]. Spearman’s rank correlation coefficient [3] has also found new use in this area (for example in [18, 19, 20]). Finally, the most recent method we examine is AUSE [2], used in for example [21, 22, 23]. While other uncertainty assessment methods exist, we have chosen to focus on these four as they are most commonly used.
III Theory
@@
First, we introduce different types of uncertainty in Sec. III-A. Then, in Sec. III-B, we introduce the four common uncertainty evaluation metrics to be analyzed.
III-A Different types of Uncertainty
@@
In the realm of predictive uncertainty estimation, we encounter two fundamental types of uncertainty [24]. The first, known as aleatoric uncertainty, encompasses the intrinsic noise and ambiguity present in observations. This noise, stemming from sources like sensor or motion irregularities, persists even with an increase in data collection and cannot be mitigated. The second type, epistemic uncertainty constitutes the uncertainty surrounding model parameters, reflecting our lack of knowledge about the precise model generating the observed data. This form of uncertainty tends to diminish as more data is acquired, and is thus often termed model uncertainty. The most common cause of epistemic uncertainty is out-of-distribution (OOD) data. In other words, data that comes from a distribution that is different from the training data. For example, autonomous driving models that are trained on synthetic data usually face domain gap to real-world data.
@@
Aleatoric uncertainty can be subdivided into homoscedastic uncertainty, which remains constant across various inputs, and heteroscedastic uncertainty, which varies depending on the model inputs. Heteroscedastic uncertainty plays a crucial role in computer vision tasks. For example, in depth regression, images with intricate textures and prominent vanishing lines typically yield confident predictions, whereas images of featureless surfaces are expected to result in significantly higher uncertainty.
III-B Uncertainty Evaluation Metrics
@@
Here, we introduce the four selected uncertainty metrics for the deep regression tasks. We define the metrics as functions of a dataset $S=\{(\mathbf{x}_{i},y_{i})\ |\ i=1,2,\ldots,N\}$, consisting of input-output pairs $(\mathbf{x}_{i},y_{i})$.
For any learned model, we further define $F(\mathbf{x};\theta)\mapsto\mathbb{R},\ \ \text{and}\ \ U(\mathbf{x};\theta)\mapsto\mathbb{R}\enskip$ to be functions based on the model parameters $\theta$. $F(\mathbf{x};\theta)$ produces a prediction of $y$, and $U(\mathbf{x};\theta)$ produces an uncertainty estimate.
AUSE: Sparsification plots are widely used in optical flow [25, 26] and stereo disparity [27] to assess how well the predicted uncertainty coincides with the prediction error on a test dataset. Ilg et al.[2] define in natural language the numerical score AUSE for the comparison among approaches with different main task performances. Here we present a formal mathematical definition.
@@
Based on a parameter $\alpha\in[0,1]$, we partition the set $S$ into disjoint subsets $S^{\operatorname{AE}}_{\land}(\alpha)$ and $S^{\operatorname{AE}}_{\lor}(\alpha)$ such that $|S^{\operatorname{AE}}_{\land}(\alpha)|=\alpha|S|$, and
@@
With these partitions in place, we define the area under sparsification error (AUSE):
$\operatorname{AUSE}(S)=$
$\frac{1}{\operatorname{MAE}(S)}\int_{0}^{1}\operatorname{MAE}(S^{U}_{\lor}(\alpha))-\operatorname{MAE}(S^{\operatorname{AE}}_{\lor}(\alpha))d\alpha.$ (4)
AUSE computes the normalized area between the two curves $\operatorname{MAE}(S_{\lor}^{U}(\alpha))$ and $\operatorname{MAE}(S_{\lor}^{\operatorname{AE}}(\alpha))$ formed by varying $\alpha$ from 0 to 1. The curve $\operatorname{MAE}(S_{\lor}^{\operatorname{AE}}(\alpha))$ is called the oracle, which represents a lower bound on the possible MAE of $S_{\lor}^{U}(\alpha)$. In practice, the integral in (III-B) is replaced with summing over $\alpha$$=$$[0,1/N,\cdots,1]$ since the test set is finite. See Fig. 5 for an example.
Calibration Error The reliability diagram [28] and the expected calibration error (ECE) [5] are originally diagnostic tools for classification models that compare sample accuracy against predicted confidence. However, this approach does not apply to regression tasks. Therefore, Kuleshov et al. [12] introduce a calibration plot for regression tasks in terms of the cumulative distribution function from the model, and summarize it with calibration error as a numerical score. Formally, let the predicted probability distribution of an input $\mathbf{x}_{i}$ be $P_{\theta}^{i}=F(\mathbf{x}_{i};\theta).$ We have the value of the predicted cumulative distribution function for the ground truth $y_{i}$ as
$P(y_{i}|\theta)=P_{\theta}^{i}(y\leq y_{i}|\theta).$ (5)
The empirical frequency is then defined as
$\hat{p}_{j}=\frac{|\{y_{i}\mid P(y_{i}|\theta)\leq p_{j},(\mathbf{x}_{i},y_{i})\in S\}|}{N}\enskip.$ (6)
@@
Spearman correlation In 1904, psychologist Charles Spearman defined what he called the rank method of correlation [3]. He observed that there may exist correlations that would not be adequately captured by simple linear correlation. Instead, he argued that more complex correlations may be adequately captured by comparing ranks of elements.
Given a set of samples, the rank of a sample is simply the index that number would have in a list. More formally, let $L=[x_{1},x_{2},\ldots,x_{N}]$ be a sequence of samples, then the rank of a sample is defined as:
$r(x_{i})=1+|\{x_{j}\ |\ x_{j}<x_{i},\ x_{j}\in L\}|\enskip.$ (8)
@@
When used as a metric for uncertainty, this correlation is computed between the predictive uncertainty, and the absolute prediction errors.
NLL The popular negative log-likelihood is proved to be a strictly proper scoring rule [4]. Given a dataset $S=\{(\mathbf{x}_{i},y_{i})\ |\ i=1,2,\ldots,N\}$ and a probability density function $p(y|\mathbf{x})$ parameterised by learnable parameters $\theta$, the NLL is defined as
@@
In expectation, the NLL is minimized if and only if $p(y_{i}\ |\ \mathbf{x}_{i};\theta)$ is equal to the true underlying data distribution [29]. As such, the NLL can also be used as a metric for uncertainty predictions, since a model with a lower NLL does a better job (in expectation) of fitting the true data distribution.
III-C Regression Models with Uncertainty Predictions
There exist many different model architectures that incorporate uncertainty predictions. However, our goal is not to investigate properties in the models. We choose to only use two different models, namely an ensemble and an energy-based model.
III-C1 Deep Ensemble (DE)
Ensemble learning combines the predictions from multiple individual models to achieve better predictive performance than any single model. For estimating predictive uncertainty, Deep Ensemble (DE) [4] is a simple yet effective method that trains multiple models in the same architecture with different random initialization and data shuffling. The ensemble is treated as a uniformly weighted mixture model. In practice, for the regression task, each individual model outputs two scalars, which are interpreted as the mean and variance of a Gaussian distribution, and then it is trained to minimize the Gaussian NLL loss as in (12) on the training set.
@@ III-C1 Deep Ensemble (DE)
We follow the suggestion in [4] to train 5 models and thus get a mixture of Gaussians. This mixture is further approximated by a Gaussian whose mean and variance are the mean and variance of the mixture respectively.
III-C2 Energy Based Regression (EBR)
Energy-based learning involves learning an energy function $\mathcal{E}(x)$. The goal when learning $\mathcal{E}(x)$ is to assign low energy to observed samples [30]. Learning $\mathcal{E}(x)$ is in many ways analogous to learning a probability density function, and as such, it has commonly been used for unsupervised learning tasks [30]. Gustafsson et al. [30] construct a supervised regression model based on an energy function. First, they define $f_{\theta}(x,y)\mapsto\mathbb{R}$ to be their learned energy function. Then, they explicitly construct a probability density function:

### v4

@@ Abstract
When deploying deep neural networks on robots or other physical systems, the learned model should reliably quantify predictive uncertainty. A reliable uncertainty allows downstream modules to reason about the safety of its actions. In this work, we address metrics for uncertainty quantification. Specifically, we focus on regression tasks, and investigate Area Under Sparsification Error (AUSE), Calibration Error (CE), Spearman’s Rank Correlation, and Negative Log-Likelihood (NLL). Using multiple datasets, we look into how those metrics behave under four typical types of uncertainty, their stability regarding the size of the test set, and reveal their strengths and weaknesses. Our results indicate that Calibration Error is the most stable and interpretable metric, but AUSE and NLL also have their respective use cases. We discourage the usage of Spearman’s Rank Correlation for evaluating uncertainties and recommend replacing it with AUSE.
journal: Pattern Recognition Letters
https://doi.org/10.1016/j.patrec.2024.09.011
[…]
[1] organization=Lund University LTH, city=Lund, country=Sweden,
[…]
[2] organization=Linköping University, city=Linköping, country=Sweden,
1 Introduction
@@
In recent years, there has been a rapid advance in the adoption of neural network-based methods in many real-world applications, for example robotics. Following this adoption, increasing scrutiny has been directed towards neural network-based methods for their lack of reliability and interpretability. While neural networks have achieved impressive performance for many different tasks, the fact remains that they can be unreliable in real-world deployment (Grimmett et al., 2016). Additionally, their lack of interpretability makes it difficult to know how and when they may perform unreliably. For these reasons, increasing attention has been directed at the uncertainty output from neural networks, and the importance of introspective qualities (Grimmett et al., 2016). Arguably the most important introspective quality is a reliable uncertainty estimate.
@@
Figure 1: An illustration for UQ metrics and regression metrics. Note: the axes of CE and AUSE are distinct, but not orthogonal.
Despite increasing attention being directed toward uncertainty quantification (UQ), much of this work focuses on uncertainty in classification tasks where metrics for assessment are well understood (Ovadia et al., 2019). Many real-world applications, however, rely instead on regression, and there is a lack of common understanding surrounding the available metrics for regression. In this work, we have identified four metrics that are commonly used to measure various qualities in a predicted uncertainty in regression. Specifically, we investigate the Area Under Sparsification Error (AUSE) (Ilg et al., 2018), Spearman’s Rank Correlation (Spearman, 1904), Negative Log Likelihood (NLL)(Lakshminarayanan et al., 2017), and Calibration Error (CE) (Pakdaman Naeini et al., 2015). These UQ metrics measure different aspects of uncertainty that are all orthogonal to the regression task performance (typically mean squared error (MSE)) as indicated in Fig. 1.
@@
With the help of synthetic datasets, we explore the usage of these metrics with the aim of better understanding their differences, strengths and limitations.
@@
  • 1.
    We create simple synthetic datasets that highlight different types of uncertainty, and use these to explore how the metrics behave, compared to the generating data distribution.
  • 2.
@@
  • 3.
    We reason about the strengths and limitations of each metric.
  • 4.
    Finally, we experiment on a real stereo disparity task, where we add synthetic noise to the input, to further illustrate the behaviour of the tested measures.
@@
We also provide a mathematical formulation for the AUSE metric. To the best of our knowledge, AUSE has previously only been described informally, using natural language.
2 Related Work
@@
Historically, methods that deal with any type of uncertainty in regression have been formulated using parameterized distributions. As a classic example, consider Kalman filters (Kalman, 1960), which are built on a parameterized distribution with well-defined uncertainty. Modern machine learning methods are built upon a solid statistical foundation and most methods output some form of uncertainty (Bishop, 2006). However, it has been observed that there is a discrepancy between the predicted uncertainty from modern neural networks, and the observed empirical accuracy (Guo et al., 2017). Additionally, lacking interpretability in these uncertainties have led to the development of methods to formulate more statistically grounded uncertainties (Gawlikowski et al., 2023).
@@
Along with the research on uncertainty estimation, a number of methods have been developed to assess how trustworthy a predicted uncertainty is. Perhaps the most commonly used (for example in Guo et al. (2017); Kuleshov et al. (2018)) is the expected calibration error (ECE) (Pakdaman Naeini et al., 2015). While not as common as ECE, NLL (negative log likelihood) is also used, for example in Heiss et al. (2022); Loquercio et al. (2020). Spearman’s rank correlation coefficient (Spearman, 1904) has also found new use in this area (for example in Tan et al. (2023); Ng et al. (2023)). Finally, the most recent method we examine is AUSE (Ilg et al., 2018), used in for example Eldesokey et al. (2020); Xiong et al. (2024). While other uncertainty assessment methods exist, we have chosen to focus on these four as they are most commonly used.
3 Theory
@@
First, we introduce different types of uncertainty in Sec. 3.1. Then, in Sec. 3.2, we introduce the four common uncertainty evaluation metrics to be analyzed.
3.1 Different types of Uncertainty
@@
In the realm of predictive uncertainty estimation, we encounter two fundamental types of uncertainty (Kendall & Gal, 2017). The first, known as aleatoric uncertainty, encompasses the intrinsic noise and ambiguity present in observations. This noise, stemming from sources like sensor or motion irregularities, persists even with an increase in data collection and cannot be mitigated. The second type, epistemic uncertainty constitutes the uncertainty surrounding model parameters, reflecting our lack of knowledge about the precise model generating the observed data. This form of uncertainty tends to diminish as more data is acquired, and is thus often termed model uncertainty. The most common cause of epistemic uncertainty is out-of-distribution (OOD) data. In other words, data that comes from a distribution that is different from the training data. For example, autonomous driving models that are trained on synthetic data usually face domain gap to real-world data.
@@
Aleatoric uncertainty can be subdivided into homoscedastic uncertainty, which remains constant across various inputs, and heteroscedastic uncertainty, which varies depending on the model inputs. Heteroscedastic uncertainty plays a crucial role in computer vision tasks. For example, in depth regression, images with intricate textures and prominent vanishing lines typically yield confident predictions, whereas images of featureless surfaces are expected to produce higher uncertainty.
3.2 Uncertainty Evaluation Metrics
@@
Here, we introduce the four selected uncertainty metrics for the deep regression tasks. We define the metrics as functions of a dataset $S=\{(\mathbf{x}_{i},y_{i})\ |\ i=1,2,\ldots,N\}$, consisting of input-output pairs $(\mathbf{x}_{i},y_{i})$. For any learned model, we further define $F(\mathbf{x};\theta)\mapsto\mathbb{R},\ \ \text{and}\ \ U(\mathbf{x};\theta)\mapsto\mathbb{R}\enskip$ to be functions based on the model parameters $\theta$. $F(\mathbf{x};\theta)$ produces a prediction of $y$, and $U(\mathbf{x};\theta)$ produces an uncertainty estimate.
AUSE: Sparsification plots are widely used in optical flow (Mac Aodha et al., 2013; Wannenwetsch et al., 2017) and stereo disparity (Häger et al., 2021) to assess how well the predicted uncertainty coincides with the prediction error on a test dataset. A sparsification plot is created by plotting the mean absolute error for a dataset while iteratively removing the data point with the highest predicted uncertainty. If the predicted uncertainty coincides well with the error, the sparsification plot should form a monotonically decreasing curve. The best possible sparsification plot, commonly called oracle sparsification, is generated when the predicted uncertainty corresponds exactly to the absolute error. Ilg et al. (2018) define sparsification error as the difference between a models’ sparsification and the corresponding oracle. AUSE is then defined as the area under the sparsification error curve. However, Ilg et al. (2018) never provide a formal mathematical definition for AUSE, and therefore we provide one here.
@@
Based on a parameter $\alpha\in[0,1]$, we partition $S$ into disjoint subsets $S^{\operatorname{AE}}_{\land}(\alpha)$ and $S^{\operatorname{AE}}_{\lor}(\alpha)$ such that $|S^{\operatorname{AE}}_{\land}(\alpha)|=\alpha|S|$, and
@@
With these partitions in place, we define:
$\operatorname{AUSE}(S)=\int_{0}^{1}\frac{\operatorname{MAE}(S^{U}_{\lor}(\alpha))}{\operatorname{MAE}(S)}-\frac{\operatorname{MAE}(S^{\operatorname{AE}}_{\lor}(\alpha))}{\operatorname{MAE}(S)}\enskip d\alpha.$ (4)
AUSE computes the normalized area between the two curves $\operatorname{MAE}(S_{\lor}^{U}(\alpha))$ and $\operatorname{MAE}(S_{\lor}^{\operatorname{AE}}(\alpha))$ formed by varying $\alpha$ from 0 to 1. The curve $\operatorname{MAE}(S_{\lor}^{\operatorname{AE}}(\alpha))$ is called the oracle, which represents a lower bound on sparsification. In practice, the integral (4) is replaced with summing over finite test set samples. See Fig. 5 for an example of an AUSE sparsification plot.
Since AUSE is computed as the difference between two curve, the definitive lower bound is zero, but there is no clear upper bound.
Calibration Error The reliability diagram (Niculescu-Mizil & Caruana, 2005) and the expected calibration error (Pakdaman Naeini et al., 2015) are originally diagnostic tools for classification models that compare sample accuracy against predicted confidence. However, this approach does not apply to regression tasks. Therefore, Kuleshov et al. (2018) introduce a calibration plot for regression tasks in terms of the cumulative distribution function from the model, and summarize it with CE as a numerical score. Formally, let the predicted probability distribution of an input $\mathbf{x}_{i}$ be $P(y_{i}|\theta)=F(\mathbf{x}_{i};\theta).$ The empirical frequency is then defined as
$\hat{p}_{j}=\frac{|\{y_{i}\mid P(y<y_{i}|\theta)\leq p_{j},(\mathbf{x}_{i},y_{i})\in S\}|}{N}\enskip.$ (5)
@@
Spearman correlation In 1904, psychologist Charles Spearman defined what he called the rank method of correlation (Spearman, 1904). He observed that there may exist correlations that would not be adequately captured by simple linear correlation. Instead, he argued that more complex correlations may be adequately captured by comparing ranks of elements.
Given a set of samples, the rank of a sample is simply the index that number would have in a list. More formally, let $L=[u_{1},u_{2},\ldots,u_{N}]$ be a sequence of samples, then the rank of a sample is defined as:
$r(u_{i})=1+|\{u_{j}\ |\ u_{j}<u_{i},\ u_{j}\in L\}|\enskip.$ (7)
@@
When used as a metric for uncertainty, this correlation is computed between the predictive uncertainty, and the absolute prediction errors (Tan et al., 2023).
NLL The popular negative log-likelihood is proved to be a strictly proper scoring rule (Lakshminarayanan et al., 2017). Given a dataset $S=\{(\mathbf{x}_{i},y_{i})\ |\ i=1,2,\ldots,N\}$ and a probability density function $p(y|\mathbf{x})$ parameterised by learnable parameters $\theta$, the NLL is defined as
@@
In expectation, the NLL is minimized if and only if $p(y_{i}\ |\ \mathbf{x}_{i};\theta)$ is equal to the true underlying data distribution (Hastie et al., 2009). As such, the NLL can also be used as a metric for uncertainty predictions, since a model with a lower NLL does a better job (in expectation) of fitting the true data distribution.
3.3 Regression Models with Uncertainty Predictions
There exist many different model architectures that incorporate uncertainty predictions. Note that our goal is not to investigate properties in the models. We choose to only use two different models, namely an ensemble and an energy-based model. While there are many other popular models for UQ (e.g. the recently introduced SNGP by Liu et al. (2024)) the purpose of this paper is to evaluate the metrics, not the models.
3.3.1 Deep Ensemble (DE)
Ensemble learning combines the predictions from multiple individual models to achieve better predictive performance than any single model. For estimating predictive uncertainty, Deep Ensemble (DE) (Lakshminarayanan et al., 2017) is a simple yet effective method that trains multiple models in the same architecture with different random initialization and data shuffling. The ensemble is treated as a uniformly weighted mixture model. In practice, for the regression task, each individual model outputs two scalars, which are interpreted as the mean and variance of a Gaussian distribution, and then it is trained to minimize the Gaussian NLL loss as in (11) on the training set.
@@ 3.3.1 Deep Ensemble (DE)
We follow the suggestion by Lakshminarayanan et al. (2017) to train 5 models and thus get a mixture of Gaussians. This mixture is further approximated by a Gaussian whose mean and variance are the mean and variance of the mixture respectively.
3.3.2 Energy Based Regression (EBR)
Energy-based learning involves learning an energy function $\mathcal{E}(x)$. The goal when learning $\mathcal{E}(x)$ is to assign low energy to observed samples (Gustafsson et al., 2019). Learning $\mathcal{E}(x)$ is in many ways analogous to learning a probability density function, and as such, it has commonly been used for unsupervised learning tasks (Gustafsson et al., 2019). Gustafsson et al. (2019) construct a supervised regression model based on an energy function. First, they define $f_{\theta}(x,y)\mapsto\mathbb{R}$ to be their learned energy function. Then, they construct a probability density function: